# Online Multivariate Changepoint Detection: Leveraging Links With Computational Geometry

### v2

@@ 5.1 Convex hull for CUSUM e-detectors
Following equations (27) of [37] and considering an exponential baseline increment as in their equation (41) we can write their update of the e-detector CUSUM statistics for a given $\lambda$ as
@@
where $\lambda$ is in a subset of $\mathbb{R}$, $s$ and $v$ are real-valued functions, and $\psi$ is a finite and strictly convex function. Taking the $\log$ and considering all possible changepoints prior to $n$ we recover with induction that
@@
By subtracting $\sum_{t=1}^{n}(s(x_{t})\lambda-\psi(\lambda)v(x_{t}))$ in the previous bracket, we do not change the maximum of the argument, and we recover our functional pruning framework of Equation (9) taking for any changepoint at $\tau<n$, $a_{\tau}=\sum_{t=1}^{\tau}v(x_{t})$, $2b_{\tau}=\sum_{t=1}^{\tau}s(x_{t}),$ and $c_{\tau}=0$. Applying Theorem 4 and its proof we get that the changepoint optimising the e-detector CUSUM score for a given $\lambda$ value corresponds to a vertex on the convex hull of the following points in $\mathbb{R}^{2}$:
@@
As argued in [37] it makes sense to consider a mixture of e-detectors, that is different $\lambda$ values. For computational reasons [37] considered a finite number of values. In practice, this number is determined using their computeBaseLine function that is detailed in their appendix B in Algorithm 3.
@@ 5.2 A parametric approach to the Plus-Minus score
Considering a change in the mean of the Gaussian model to fit the Plus-Minus score of the Cavaliers seems rather naive, as it does not take into account the discrete nature of the data and possible dependencies. In this context, we argue that it is also not clear how one could get a valid estimation of the pre-changepoint parameter. Finally, as argued in [37] and shown a bit later calibrating the threshold to detect a changepoint is not trivial. Roughly speaking the threshold depends on the choice of the null distribution and this largely affects the position at which a change is detected. Here as a robust parametric solution to these problems, we consider a change in the mean and variance of a univariate Gaussian signal with unknown pre-changepoint parameters.
For a change in mean and variance using our convex hull framework Equation (9) and following its natural parametrisation presented in Section 2.2 we need to keep track of the vertices on the hull of a 3-dimensional set of points
@@
As we will see this second model, although still parametric, is much more robust to the choice of the null distribution.
Interestingly, this mean and variance model is invariant to shift and scaling. Based on that, we see that all the points on the convex hull of the Plus-Minus e-detector are on the hull of the change in mean and variance model. Indeed for this second e-detector the raw Plus-Minus score was normalized as follows $x^{\prime}_{t}=(x_{t}+80)/160$ and then $s$ and $v$ were chosen to be $s(x^{\prime})=(x^{\prime}/m-1)$ and $v(x^{\prime})=s(x^{\prime})^{2})$. Therefore all changepoints tested by the Plus-Minus e-detector will also be tested by our mean and variance model.
As the data is discrete it is not unlikely that two consecutive scores are exactly equal (this happens 19 times out of 639 in the Plus-Minus score of the Cavaliers from 2010 to 2018). In fact, we also observed 2 stretches of 3 equal scores. For such stretches the estimated variance would be 00 and the likelihood would blow up to infinity. To get around this problem we considered a minimum value for the variance. We computed the per-season variance of all NBA teams from season 2003-2004 to season 2018-2019 and observed a minimum of $97.02$, a maximum of $247.38$ and a mean of $170.68$. Hence, we argue that setting a minimum variance of 1 (which is roughly a hundred times smaller) should not hamper our ability to detect changes, while also limiting the impact of small segments with equal values.
@@
Figure 7: CUSUM statistics for the mean model for the mean and variance model. The dotted line corresponds to the beginning of season 2014-15 and match number 312.
Interestingly, at any time-point after 320 (8th match of season 2014-15) the maximum likelihood changepoint is always 279 (5th of February 2014) for both models (mean and, mean and variance). Furthermore, the convex hull for the mean and variance model (in Equation (13)) - and hence the mean model and Plus-Minus e-detector (recall that all changepoints tested by the Plus-Minus e-detector are tested by the mean and variance model) - no longer contains vertex 312 after time $n=320$ (8th match of season 2014-15).
Match 279 was played on the 5th of February 2014, well before the beginning of season 2014-15 in October 2014. At first, this might come as a surprise. But, it indeed seems that the mean Plus-Minus score increased after the 6th of February 2014 (average score of $-6.28$ before compared to $1.12$ after and significant t-test p-values of $0.01584$). See also the zoom on the Plus-Minus score over season 2013-14 and 2014-15 in Figure 8. Also, as discussed on the Wikipedia web page of David Griffin [70] “On February 6, 2014, he was named the acting general manager for the Cavaliers” and David Griffin, was the one ”to sign LeBron James back” for season 2014-15.
@@
Figure 8: Plus-Minus score of the Cavaliers from seasons 2013-14 to 2014-15 as a bar plot. Losses are in reddark red and wins are in dark blue. The ends of seasons are represented with vertical dashed lines (grey) and seasons are shown on the bottom with grey arrows.
5.2.1 Simulating NBA Plus-Minus scores without changepoint
@@
Based on Figure 7 it is clear that for a particular range of the threshold both the mean and mean and variance model with unknown pre-changepoint parameter can detect a changepoint after the beginning of season 2014-15. To calibrate this threshold one typically needs a simulation model without a changepoint. Here, we will set the CUSUM threshold such that under our null simulation over 1000 matches we would detect a change only in 5% of the cases.
Our idea here is not to precisely model the Plus-Minus score of the Cavaliers but rather explore a few natural optimistic and pessimistic options to model it and check how this affects the threshold of the two models and in turns the time at which they detect the change.
The Plus-Minus score is integer-valued. A straightforward, albeit maybe simple, choice would be to consider the difference of two independent Poisson distributions. This is the so-called Skellam distribution. It has two parameters (the mean of the first and second Poisson $\lambda_{1}$ and $\lambda_{2}$). We simulated data using values of $\lambda_{1}=\lambda_{2}$ both equal to 95, 105, and 115. We considered mean values around $100$ as the points per team of an NBA match from season 1999-00 to season 2021-22 range from $93.40$ to $112.09$ with a mean of $101.07$.
A slightly more complex choice is to consider the difference of two independent Negative Binomial distributions with equal scale parameters, $r_{1}$ and $r_{2}$. As for the Poisson we considered $\lambda_{1}=\lambda_{2}$, which were set to $95$, $105$, and $115$, along with scale parameters $r_{1}=r_{2}$ at values of $500$, $100$, $50$, and $10$. This makes a total of $12$ combinations.
A data-driven way to model the Plus-Minus score without changepoint is to re-sample past games. We did this sampling uniformly at random past matches of the Cavaliers from season 1999-2000 to 2009-2010 or uniformly at random past matches of any NBA team from season 1999-2000 to 2009-2010. Finally, following the argument of [37] and noting that the Plus-Minus score is never larger than $80$ we simulated integer valued score uniformly at random between $-80$ and $80$.
In the Poisson and Negative Binomial based model, we could further model dependencies in the data. To do this we varied the mean of the first Poisson (or Negative Binomial) over time. We considered two scenarios. In the first, we added to the mean an auto-regressive process of order 1 and coefficient $0.6$ multiplied by either 1, 2, 4, or 8. In the second we added sin waves with a period of 20 matches multiplied by either 1, 2, 4, or 8. We present in Figure 9 two simulated score profiles.
(a) Using AR(1) with AR parameter $0.6$ (multiplied by $8$)
(b) Using sin waves with a period of 20 matches (multiplied by $8$)
@@
Figure 9: Two Plus-Minus score simulation examples based on the Negative Binomial model with a mean of $115$ and scale parameter of $50$ using two dependency scenarios.
Figure 10(a) represents the various thresholds (in log-scale) we get for our two models, mean and mean and variance. The corresponding detection times are represented in Figure 10(b).
@@
It can be noted that the threshold is much more variable for the changepoint in the mean model and this translates to a much more variable detection time. To be specific for the mean model, the change is detected at the earliest at match 330 (the 5th of December 2014), but in $21.7\%$ of the cases the threshold is so high that the changepoint is never detected (reported in the histogram as a detection at $1000$). For the mean and variance model, the change is always detected at the earliest at match 339 (the 23rd of December 2014) and at the latest at match 368 (the 20th of February 2015). In fact in $94.2\%$ of the simulation models the change is detected before with the mean and variance model.
Based on this numerical application we argue that being able to run efficiently a bi-parametric model online can be beneficial even if the model (Gaussian here) does not properly fit the nature (discrete here) of the data.
(a) Threshold
(b) Detection time (1000 = not detected)
@@
Figure 10: Threshold and Detection time obtained for the mean and mean and variance model.
@@ 6 Discussion
In this paper, we establish a connection between the detection of a single changepoint in $p$-dimensional data streams and a convex hull problem in dimension $p+1$. Assuming independence of all observations and dimensions and a continuous distribution we demonstrate using the recent probabilistic paper [57] that the expected number of faces and vertices of the convex hull is of order $2(\log(n))^{p}$ where $n$ is the number of data points.
Based on this, we derive a simple yet efficient algorithm, MdFOCuS Sp, that allows maximising the full or restricted likelihood-ratio test statistics (that is allowing only $s$ coordinates to change with $s\leq p$) over all possible window sizes. MdFOCuS Sp generalises the FOCuS algorithm [34] to the multivariate case. We show that MdFOCuS Sp is empirically fast for up to $p=5$ and allows to process more than $10^{5}$ data points in a matter of minutes. For larger $p$ we provide an approximation storing only $\mathcal{O}(p\log^{\tilde{p}}(n))$ vertices of the convex hull where $\tilde{p}\leq p.$ For $\tilde{p}=2$ this approximation is suitable for high-dimensional data. We also show using simulation that MdFOCuS Sp and its approximation (MdFOCuS $\tilde{p}$d Sp) are statistically competitive in particular when the change affects all dimensions or when the parameter prior to the change is unknown.
A few improvements to the computational complexity of the procedure could be made to allow the procedure to scale to higher dimensions and longer sequences. The most computationally expensive components of the procedure are respectively the QuickHull algorithm, for pruning the list of indices, and the maximisation step. We mentioned already how further research could go towards a sequential update of the boundary of the convex hull of our points, to improve on the complexity of QuickHull. The idea would be to employ information from the hull at the previous iteration, evaluate the inclusion of the most recent point, and update the faces of the neighbouring vertex consequently.
Recently, [71] offered a solution to improve on the maximisation step. Instead of looking at the maxima of all indices to calculate the test statistic, the authors exploit a bound on the maxima of the differences of contiguous indices. The main idea is that the overall maximum, our test statistics, is bounded by the sum of those maxima of the differences. If this sum does not pass the threshold there is no need to compute the likelihood ratio-statistics for all unpruned changepoints. Fast sequential updates of the sum of the maxima of the difference exists, effectively rendering the univariate algorithm constant per iteration [71]. Whilst this bound translates exactly to the multivariate case, it is easily much larger than the threshold, given that we need to store and sum a large amount of indices even in dimension 2. A less conservative bound could be the object of future study to improve on the maximisation step in higher dimensions.
To conclude, our mathematical argument is valid for a large collection of parametric online single changepoint models with distributions taken from the exponential family (or more generally with a likelihood that can be algebraically written as the natural exponential family). Note that this includes models where data are of different types (e.g. with $p=2$ a first Poisson component, and a second Gaussian component). We also show in Supplementary Material D that there is a connection between multiple changepoint models and the half-space intersection problem of Theorem 3. Furthermore, we show in Section 5 that Theorem 5 could be beneficial to speed the calculation of the non-parametric e-detectors. Finally, note that no assumptions are made on the set of indices $\mathcal{T}$ in Equation (9) and for example, our theory also applies to changepoint models with observations on a tree rather than a chain [72]. For all these reasons we argue that this connection between changepoints and convex hull (or half-space intersection) sheds light on the geometric nature of the changepoint detection problem and is a promising line for future research.
Supplementary Material
@@
Appendix A Proofs
A.1 The proof of Theorem 2
Proof.
Consider that $\tau$ is in $\mathcal{G}_{\mathcal{T}}$. By definition, there exists $(\lambda_{0},\mu_{0})\in Im(A)\times\mathcal{D}_{A}=\mathbb{R}^{+}\times\mathbb{R}^{p}$ such that for all $\tau^{\prime}\neq\tau$ in $\mathcal{T}$ we have: $g_{\tau}(\lambda_{0},\mu_{0})>g_{\tau^{\prime}}(\lambda_{0},\mu_{0})$. Using condition $c_{\tau}=c_{\tau^{\prime}}=0$ and multiplying by $\alpha>0$ we get by linearity
@@ A.1 The proof of Theorem 2
leading to solution $\alpha_{0}=\frac{\lambda_{0}^{\frac{1}{q-1}}}{\|\mu_{0}\|\frac{q}{q-1}}$. $\alpha_{0}$ is well defined as we excluded $\|\mu_{0}\|=0$ and $\lambda_{0}=0$. Further, note that the result is still true for $a_{\tau}=0$ or $a_{\tau^{\prime}}=0$. ∎
A.2 The proof of Theorem 4
Proof.
@@ A.2 The proof of Theorem 4
To go from (2) to (3) we note that a point is a part of the convex hull if there is a hyperplane (defined by a non-zero vector and a constant) going through this point and such that all other points are strictly on one side of the hyperplane. ∎
A.3 The proof of Theorem 5
Proof.
First consider the case $s=p.$ Starting from equation (6), we get to equation (8) as described in section 2.2 and we recover our functional pruning problem of equation (10) as described in Section 3.1, taking $\mathcal{T}=\{1,\ldots n-1\}$, $a_{\tau}=\tau$, $b_{\tau}=\sum_{t=1}^{\tau}x_{t}$ and $c_{\tau}=0$. Thus applying Theorem 4, and then Theorem 1 we get the desired result for $\hat{\tau}(\eta_{1})$.
@@ A.3 The proof of Theorem 5
Without loss of generality let us now consider the $s$ first coordinates of $x_{t}.$ We call $x^{\prime}_{t}$ the vector of $\mathbb{R}^{s}$ containing the first $s$ coordinates of $x_{t}$ and define $P^{\prime}(\tau)=\left(\tau,\sum_{t=1}^{\tau}x^{\prime}_{t}\right).$ Now using the previous result the change maximizing the likelihood restricted to the first $s$ coordinates is on the hull of $P^{\prime}(\tau)$ and therefore also on the hull of $P(\tau)$. ∎
A.4 The proof of Lemma 1

### v3

@@ 5.1 Convex hull for CUSUM e-detectors
Following Equations (27) and (41) of Shin et al. (2022) we can write their update of the e-detector CUSUM statistics for a given $\lambda$ as
@@
where $\lambda$ is in a subset of $\mathbb{R}$, $s$ and $v$ are real-valued functions, and $\psi$ is a finite and strictly convex function. Taking logs and considering all possible changepoints prior to $n$ we recover with induction that
@@
By subtracting $\sum_{t=1}^{n}(s(x_{t})\lambda-\psi(\lambda)v(x_{t}))$ in the previous bracket, we do not change the maximiser, and we recover our functional pruning framework of Equation (8) with $a_{\tau}=\sum_{t=1}^{\tau}v(x_{t})$, and $2b_{\tau}=\sum_{t=1}^{\tau}s(x_{t})$. Applying Theorem 4 we get that the changepoint optimising the e-detector CUSUM score for any $\lambda$ value corresponds to a vertex on the convex hull of
@@
As argued in Shin et al. (2022) it makes sense to consider a mixture of e-detectors, that is different $\lambda$ values. For computational reasons Shin et al. (2022) considered a finite number of values, chosen using their computeBaseLine function.
@@ 5.2 A parametric approach to the Plus-Minus score
Considering a change in the mean of a Gaussian model to fit the Plus-Minus score of the Cavaliers does not take into account the discrete nature of the data and possible dependencies. It is also not clear how one could get a valid estimation of the pre-changepoint parameter. Finally, as argued in Shin et al. (2022) calibrating the threshold to detect a changepoint is not trivial. As a robust parametric solution to these problems, we consider a change in the mean and variance of a univariate Gaussian signal with unknown pre-changepoint parameters.
For a change in mean and variance using our convex hull framework (8) we need to keep track of the vertices on the hull of a 3-dimensional set of points
@@
Interestingly, this mean and variance model is invariant to shift and scaling. Based on that, we see that all the points on the convex hull of the Plus-Minus e-detector are on the hull of the change in mean and variance model.
As the data is discrete, two (or more) consecutive scores may be exactly equal. For segments containing equal scores, the estimated variance would be 00, and the likelihood would blow up to infinity. To get around this problem we considered a minimum value for the variance, which we set to $1$ (chosen as substantially below the variance seen for plus-minus statistics for NBA teams).
@@
Figure 6: CUSUM statistics for the mean model (left) for the mean and variance model (right). The dotted line corresponds to the beginning of season 2014-15 and match number 312.
Interestingly, at any time-point after 320 (8th match of season 2014-15) the maximum likelihood changepoint is always between matches 279 and 280 (5th and 7th of February 2014) for both models. This is well before the beginning of season 2014-15 and the comeback of Lebron James. However, this matches the arrival of David Griffin as general manager on the 6th of February Wikipedia (2023) .
@@
Figure 7: Plus-Minus score of the Cavaliers from seasons 2013-14 to 2014-15 as a bar plot. Losses are in dark red and wins are in dark blue. The ends of seasons are represented with vertical dashed lines (grey) and seasons are shown on the bottom with grey arrows.
5.3 Simulating NBA Plus-Minus scores without changepoint
@@
Based on Figure 6 it is clear that for a particular range of the threshold both the mean and mean and variance model with unknown pre-changepoint parameter can detect a changepoint after the beginning of season 2014-15. To calibrate this threshold one typically needs a simulation model. Below we consider different choices of simulation model to see how sensitive the threshold is to this choice for both models. In all cases we will set the CUSUM threshold such that under our null simulation over 1000 matches we would detect a change only in 5% of the cases.
The simulation models we considered are: (i) the difference of two independent Poisson random variables; (ii) the difference of two negative binomial random variables; (iii) a uniform distribution on $\{-80,\ldots,80\}$ (as suggested in Shin et al., 2022); (iv) an independent resampling of past games. For (i) and (ii) we consider a range of parameters, with the mean of the random variables set to be 95, 105 or 115; and with the negative-binomial scale parameter set to 1 or 2. For (iv) we considered either sampling past matches of the Cavaliers or of NBA teams in general, from seasons 1999/2000 to 2009/10. Finally we also considered version of the scenarios in (i) and (ii) where we introduced temporal dependencies by adding to the mean a time-varying function either simulated from an AR(1) model with auto-correlation of 0.6 or using a sin function.
@@
Figure 8: Two Plus-Minus score simulation examples based on the Negative Binomial model with a mean of $115$ and scale parameter of $50$ using two dependency scenarios: an AR(1) model with AR parameter $0.6$ (left) and using sine waves with a period of 20 matches (right).
@@
Figure 9 shows the various thresholds we get for our two models, and the corresponding detection times. The threshold is much more variable for the change in mean model and this translates to a much more variable detection time. To be specific for the mean model, the change is detected at the earliest at match 330 (the 5th of December 2014), but in $21.7\%$ of the cases the threshold is so high that the changepoint is never detected (reported in the histogram as a detection at $1000$). For the mean and variance model, the change is always detected at the earliest at match 339 (the 23rd of December 2014) and at the latest at match 368 (the 20th of February 2015). In fact in $94.2\%$ of the simulation models the change is detected before with the mean and variance model.
@@
Figure 9: Threshold (left) and detection time (right) obtained for the mean model and the mean and variance model.
@@ 6 Discussion
In this paper, we establish a connection between the detection of a single changepoint in $p$-dimensional data streams and a convex hull problem in dimension $p+1$. Based on this, we derive a simple yet efficient algorithm, MdFOCuS. This algorithm finds a set of time-points that contain the points associated with all vertices of a convex hull, and that we show contains all possible times that can maximise the likelihood-ratio statistic for a change. The method is general in that in can be used for a wide-range of models from the exponential family, including multivariate data where different streams are modelled by different distributions. It can also deal with the pre-change parameter being known or unknown, and covers tests where we allow only a subset of components of a multivariate data stream change. For all these varieties, the only difference is the part of the algorithm which maximises the function associated with each time-point we have stored. We also showed that it can be used to speed up the calculation of e-detectors. MdFOCuS is efficient for $p\leq 5$. For larger $p$ we provide an approximation storing only $\mathcal{O}(p\log^{\tilde{p}}(n))$ vertices of the convex hull where $\tilde{p}\leq p.$
Improvements to the computational complexity of the procedure could be made to allow the procedure to scale to higher dimensions and longer sequences. For example, a sequential update of the boundary of the convex hull of our points, to improve on the complexity of QuickHull. Recently, Ward et al. (2024) offered an approach to speed-up the maximisation step. This aims to recycle calculations to give a bound on the test statistic that is much more efficient to calculate. Whilst this idea can directly be applied in the multivariate settings we consider, the bounds are too loose to give any noticeable speed-up. A less conservative bound could be the object of future study.
More generally, the ideas in this paper which show a connection between changepoint detection and convex hull (or half-space intersection) problems sheds light on the geometric nature of the changepoint detection problem and is a promising line for future research.
Supplementary Material for ‘Multivariate Maximum Likelihood Changepoint Detection via Half-Space Intersection and Convex Hull’
Appendix A Overview of Code
@@
Appendix B Proofs
B.1 Proof of Theorem 2
Consider that $\tau$ is in $\mathcal{G}_{\mathcal{T}}$. By definition, there exists $(\lambda_{0},\mu_{0})\in Im(A)\times\mathcal{D}_{A}=\mathbb{R}^{+}\times\mathbb{R}^{p}$ such that for all $\tau^{\prime}\neq\tau$ in $\mathcal{T}$ we have: $g_{\tau}(\lambda_{0},\mu_{0})>g_{\tau^{\prime}}(\lambda_{0},\mu_{0})$. Multiplying by $\alpha>0$ we get by linearity
@@ B.1 Proof of Theorem 2
leading to solution $\alpha_{0}=\frac{\lambda_{0}^{\frac{1}{q-1}}}{\|\mu_{0}\|\frac{q}{q-1}}$. $\alpha_{0}$ is well defined as we excluded $\|\mu_{0}\|=0$ and $\lambda_{0}=0$. Further, note that the result is still true for $a_{\tau}=0$ or $a_{\tau^{\prime}}=0$.
B.2 Proof of Theorem 4
@@
To go from point 2 to point 3 we note that a point is a part of the convex hull if there is a hyperplane (defined by a non-zero vector and a constant) going through this point and such that all other points are strictly on one side of the hyperplane.
B.3 Proof of Theorem 5
First consider the case $s=p$. Starting from (5), we get to (7) as described in Section 2.2 and we recover our functional pruning problem of equation (9) as described in Section 3.1, taking $\mathcal{T}=\{1,\ldots n-1\}$, $a_{\tau}=\tau$ and $b_{\tau}=\sum_{t=1}^{\tau}x_{t}$. Thus applying Theorem 4, and then Theorem 1 we get the desired result for $\hat{\tau}(\eta_{1})$.
@@ B.3 Proof of Theorem 5
Let us now consider the $s$ first coordinates of $x_{t}$. We call $x^{\prime}_{t}$ the vector of $\mathbb{R}^{s}$ containing the first $s$ coordinates of $x_{t}$ and define $P^{\prime}(\tau)=\left(\tau,\sum_{t=1}^{\tau}x^{\prime}_{t}\right)$. Now, using the previous result, the change maximizing the likelihood restricted to the first $s$ coordinates is on the hull of $P^{\prime}(\tau)$ and therefore also on the hull of $P(\tau)$. The same argument applies to any other choice of $s$ coordinates, and thus will apply to the set that maximises the sparse log-likelihood.
B.4 Proof of Theorem 6
The proof comes from applying results of Kabluchko et al. [2017]. To do so we need to prove that the random walk $\{P(\tau)\}_{\tau\in\{1,\ldots,n-1\}}=\{(\tau,\sum_{t=1}^{\tau}x_{t})\}_{\tau\in\{1,\ldots,n-1\}}$ satisfies two properties, $(Ex)$ and $(GP)$ defined below.
We show this in the following lemma.
Lemma 2.
Assuming that all $x_{\tau}^{d}$ are independent and identically distributed with a continuous distribution, the properties $(Ex)$ and $(GP)$ of Kabluchko et al. [2017] are satisfied.